OpenStudy (anonymous):

5-3^x=-40 round to the nearest ten thousandth i have no clue what i am doing here.

OpenStudy (phi):

is this \[ 5 -3^x = -40 \]?

OpenStudy (anonymous):

yes

OpenStudy (phi):

the first step is add -5 to both sides of the equation and then simplify: \[-5+ 5 -3^x = -5 -40 \]

OpenStudy (anonymous):

okay so you now have 3^x = -45 right

OpenStudy (phi):

almost. it is -3^x= -45 -3^x is the same as -1* 3^x so you now have \[ -1\cdot 3^x = -45 \] multiply both sides by -1

OpenStudy (anonymous):

-3^x=45

OpenStudy (phi):

multiply both sides by -1 \[ -1\cdot -1\cdot3^x = -1\cdot-45 \] what do you get ?

OpenStudy (anonymous):

3^x=45

OpenStudy (phi):

much better. now it is the same problem as the previous one.

OpenStudy (anonymous):

log 3 /log 45

OpenStudy (phi):

you are missing an important idea if you have 2*x= 4 you can get the x by itself if you divide 2x by 2 (because 2/2 is 1) and 1*x is x but if you divide the left side by 2, you must divide the right side by 2 to keep things equal. now in your problem you have log(3)*x = log(45) you want the x by itself. what do you divide by to get x by itself ?

OpenStudy (anonymous):

3

OpenStudy (phi):

if you have log(3) times x you divide by log(3). not just 3 can you finish log(3)*x = log(45)

OpenStudy (anonymous):

.2886

OpenStudy (phi):

if you have log(3) times x you divide by log(3). you have log(3)*x = log(45) what do you divide both sides by ?

OpenStudy (anonymous):

log3

OpenStudy (phi):

yes. if you divide both sides by log3 you get log(3)*x/log(3) = log(45)/log(3) can you simplify this ?

OpenStudy (anonymous):

x=log (45)/log(3)

OpenStudy (phi):

yes, and x is ?

OpenStudy (anonymous):

x=3.4650

OpenStudy (anonymous):

or is it x=.2886

OpenStudy (phi):

we can check your answer by putting this into google 5-3^(3.4650)= if should come out very close to -40

OpenStudy (phi):

the answer is not x=0.2886. It would be if we started with log(45) x = log(3) then to get "x by itself" we would divide both sides by log(45) but that is not the problem.

OpenStudy (anonymous):

-40.0013

OpenStudy (phi):

right, so in the original problem 5-3^x=-40 the left side is very close to -40 (not exact because we rounded x to the nearest ten thousandth) so you know x = 3.4650 is the correct answer.